# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v3

@@ Abstract
Combined with a Dijkstra/Prim type algorithm, this leads to a realization of the Fréchet distance (i.e., a morphing) that is locally optimal (aka locally correct), that is both easy to compute, and in practice, takes near linear time on many inputs. The new morphing has the property that the leash is always as short-as-possible. These matchings/morphings are more natural, and are better than the ones computed by standard algorithms – in particular, they handle noise more graciously. This should make the Fréchet distance more useful for real world applications.
We implemented the new algorithm, and various strategies to obtain fast practical performance. We performed extensive experiments on our new algorithm, and released publicly available (and easily installable and usable) Julia and Python packages. In particular, the Julia implementation, for computing the regular Fréchet distance, seems to be significantly faster than other currently available implementations. See Figure 2.1.
Links and information about our implementations are available here: frechet.xyz, with numerous detailed examples with animations are available here.
@@ 1.1 Definitions
Given two polygonal curves, their Fréchet distance is the length of a leash that a person needs, if they walk along one of the curves, while a dog connected by the leash walks along the other curve, assuming they synchronize their walks so as to minimize the length of this leash. (I.e. they walk so as to minimize their maximum distance apart during the walk.) Our approach is slightly different than the standard approach, and we define it carefully first.
@@ 1.1.1 Free space diagram and morphings
The free space diagram $R$ is partitioned into a non-uniform grid, where each cell corresponds to all leash lengths when a point lies on a fixed segment of one curve, and the other lies on a fixed segment of the other curve, see Figure 1.1. The sublevel set of the elevation function inside such a grid cell is a clipped ellipse.
@@
A morphing $\mathcalb{m}$ between $\pi$ and $\sigma$ is a (not self-intersecting) curve $\mathcalb{m}\subseteq R\left(\pi,\sigma\right)$ with endpoints $(0,0)$ and $(\left\|{\pi}\right\|,\left\|{\sigma}\right\|)$. The set of all morphings between $\pi$ and $\sigma$ is $\mathcal{M}_{\pi,\sigma}$. A morphing that is a segment inside each cell of the free space diagram that it visits, is well behaved [^1] All the morphings we deal with in this paper are well behaved. .
@@ 1.2 Background
Alt and Godau ag-cfdbt-95 presented a rather involved $O(n^{2}\log n)$ time algorithm to compute the Fréchet distance using parametric search. The parametric search can be removed by using randomization, giving a simpler algorithm as shown in hr-fdre-14. Buchin et al. bblmm-cfdrl-16 presented an alternative algorithm for computing the Fréchet distance that replaces the decision procedure by using a data-structure to maintain appropriate lower envelopes. Despite some simplifications, all these algorithms are somewhat involved.
Unfortunately, it is believed this problem requires quadratic time in the worst case, although a logarithmic speedup is possible; see bbmm-fswdi-17 and references therein. The quadratic time can be improved for realistic inputs by assuming that the input is “nice”, and introducing approximation, but the resulting algorithms are still not simple dhw-afdrc-12.
@@ 1.2.1 Variants of the Fréchet distance
The weak Fréchet distance allows morphings where the agents are allowed to go back and traverse portions of the curve visited previously (i.e., the morphing does not have to be $x/y$-monotone). Since the weak Fréchet distance allows considering more parameterizations, it is potentially smaller, and we have that $\mathsf{d}^{w}_{\mathcal{F}}\left(\pi,\sigma\right)\leq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, for any two curves $\pi,\sigma$. Elegantly, Alt and Godau ag-cfdbt-95 showed the weak Fréchet distance can be reduced to computing the minimum spanning tree of an appropriate graph. Unfortunately, there does not seem to be a natural way to overcome this non-monotonicity (and thus get the “strong” version).
@@
The complexity of these algorithms, together with sensitivity of the Fréchet distance to noise, lead to using “easier” related measures, such as the discrete version of the problem, and dynamic time-warping (discussed below). In the discrete version, you are given two sequences of points $p_{1},\ldots,p_{n}$, and $q_{1},\ldots,q_{m}$, and the purpose is for two “frogs” starting at $p_{1}$ and $q_{1}$, respectively, to jump through the points in the sequence until reaching $p_{n},q_{m}$, respectively, while minimizing the maximum distance between the two frogs during this traversal. At each step, only one of the frogs can jump from its current location to the next point in its sequence (no jumping back). Computing the optimal distance under this measure can be done by dynamic-programming, similar to the standard approach to edit-distance. Indeed, the configuration space here is the grid $H=\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$, where $\left\llbracket n\right\rrbracket=\{1,\ldots,n\}$.
To use the discrete version in the continuous case, one sprinkles enough points along both input curves, and then solves the discrete version of the problem. Beyond the error this introduces, to get a distance that is close to the standard Fréchet distance, one has to sample the two curves quite densely in some cases.
For the (monotone) discrete Fréchet distance the induced graph on the grid $H$ is a DAG, and the task of computing the Fréchet distance is to find a minimum bottleneck path from $(1,1)$ to $(n,m)$, where the weights are on the vertices. Here, the weight on the vertex $(i,j)$ is the distance $\left\|{p_{i}-q_{j}}\right\|$. In particular, a Fréchet morphing is an $x/y$-monotone path in $H$ from $(1,1)$ to $(n,m)$. The standard algorithm to do this traverses the grid, say, by increasing rows $i$, and in each row by increasing column $j$, such that the value at $(i,j)$ is the maximum of the length of the leash of this configuration, together with the minimum solution for $(i-1,j)$ and $(i,j-1)$. This algorithm leads to a straightforward, $O(nm)$ time algorithm for the discrete Fréchet distance. However, the morphing computed might be inferior, see Figure 1.2 for such a bad example.
@@
For simplicity, assume that the pairwise distances between all pairs of points in the two sequences are unique. We would like to imagine that we have a retractable leash, that can become shorter at times, and the leash “resists” being longer than necessary. It is thus natural to ask for a morphing where the leash as short as possible at any point in time.
Informally, the optimal retractable Fréchet morphing between the two sequences includes the bottleneck configuration, realizing the Fréchet distance, in the middle of its path, and the two subpaths from the endpoints to this configuration have to be also recursively optimal. This is formally defined and described in Section 3.2. This concept was introduced by Buchin et al. bbms-lcfm-19. Interestingly, they show that the discrete version can be computed in $O(nm)$ time, but unfortunately, the algorithm is quite complicated. They also show that the continuous retractable Fréchet can be computed in $O(n^{3}\log n)$ time.
Buchin et al. bbms-lcfm-19 refers to this Fréchet distance as locally correct, but we prefer the retractable labeling. The term “retractable Fréchet” was used by Buchin et al. bblmm-cfdrl-16, but in a different (and not formally defined) context than ours.
@@
One way to get less sensitivity for noise is to compute the total area “swept” by the leash as the walk is being performed. In the discrete case, we simply add up the lengths of the leashes during the configurations in the walk. There is also work on extending this to the continuous setting mp-cdtwt-99, bbknp-klmct-20. For the continuous case this intuitively boils down to computing (or approximating) an integral along the morphing. See Section 5.5 for more details.
@@ 1.2.2 Critical events
    Vertex-edge events: The distance between a vertex of one curve, and an edge of the other.
@@ 1.2.3 Algorithm engineering the Fréchet distance
Given the asymptotic complexity and involved implementations of the aforementioned algorithms, there has been substantial work on practical aspects of computing the Fréchet distance. In particular, in 2017, ACM SIGSPATIAL GIS Cup had a programming challenge to implement algorithms for computing the Fréchet distance. See wo-asgcr-2018 for details.
More recently, Bringmann et al. bkn-wdfpa-21 presented an optimized implementation of the decider for the Fréchet distance. Somewhat informally, Bringmann et al. bkn-wdfpa-21 builds a decider for the Fréchet distance using a $kd$-tree over the free space diagram, keeping track of the reachable regions on the boundary of each cell, refining cells by continuing down the (virtual) $kd$-tree if needed.
@@ 1.3.1 Result I: A new algorithm for retractable discrete Fréchet
We observe that a natural approach to compute the retractable Fréchet morphing is to modify Dijkstra’s/Prim’s algorithm so that it solves the minimum bottleneck path problem. This observation surprisingly seems to be new, and leads to a simpler (but log factor slower) algorithm for computing it. The only modification of Dijkstra necessary is that one always handles the cheapest edge coming out of the current cut induced by the set of vertices already handled. (In the discrete Fréchet case, the weights are on the vertices, but this is a minor issue.) This modified version of Dijkstra is well known, but we include, for the sake of completeness, the proof showing that it indeed computes the recursively optimal path, which is also a retractable Fréchet morphing between the two sequences. This immediately leads to better and more natural morphings, see Figure 1.2.
@@
Figure 1.3: Two curves, and their associated VE-Fréchet graph. Note, that every internal edge of the grid contains a portal (i.e., vertex of the graph), but in many cases the portals of two edges are co-located on their common vertex. For the edges adjacent to the starting and ending corners we set their portals to lie at the corners themselves. See here for more info.
@@ 1.3.2 Result II: A new distance and algorithm: VE-Fréchet
Because of the continuity and strict convexity of the elevation function, the function has a unique minimum on each edge of the free space diagram grid – geometrically, this is the minimum distance between a vertex of one curve, and an edge of the other curve (a vertex-edge event). We restrict our solution to enter and leave a cell only through these portals (which are easy to compute). The continuous configuration space now collapses to a discrete graph that is somewhat similar to the natural grid graph on $\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$. Indeed, a grid cell has four portals on its boundary edges. Specifically, there are directed edges from the portal on the bottom edge, to the portal on the top and right edges of the cell. Similarly, there are edges from the portal on the left edge, to the portals lying on the top and right edges. As before, the values are on the portals, and our purpose is to compute the optimal bottleneck path in this grid-like graph. Examples of this graph are depicted in Figure 1.3, Figure 4.1 and Figure 4.2.
@@
A somewhat similar idea was used by Munich and Perona mp-cdtwt-99, but they used it in the other direction – namely, in defining a better CDTW distance for two discrete sequences. However, this idea was already present (implicitly) in the original work of Alt and Godau ag-cfdbt-95 – indeed, their algorithm for the Weak Fréchet distance uses only the Vertex-Edge events (i.e., edges in the free space diagram). This boils down to solving the bottleneck shortest path problem in an undirected graph. In this case, this problem can be solved by computing the minimum spanning tree (e.g., by Prim’s algorithm, which is a variant of Dijkstra’s algorithm), as Alt and Godau do. For the directed case, one needs to use a variant of Dijkstra’s algorithm gt-atbop-88 – see Lemma 3.2. See also Buchin et al. bbdfj-ctmc-17 who also used a similar idea.
We can now run the retractable bottleneck shortest-path algorithm (i.e., the variant of Dijkstra described above) on this implicitly defined graph computing the vertices and edges of it, as they are being explored. For many natural inputs, this algorithm does not explore a large fraction of the configuration space, as it involves distances that are significantly larger than the maximum leash length needed. The algorithm seems to have near linear running time for many natural inputs. The VE-Fréchet morphing is the one induced by the computed path in this graph. Unfortunately, the VE-Fréchet morphing might allow the agents to move backwards on an edge, but importantly, the motion across a vertex is monotone. Namely, the VE-Fréchet is monotone for vertices, but not necessarily monotone on the edges. A vertex is thus a checkpoint that once passed, cannot be crossed back.
@@ 1.3.3 Result III: New algorithm for the regular Fréchet distance
One can of course turn any morphing into a monotone one by staying put instead of moving back. This is appealing for VE-Fréchet, as the corresponding VE morphing $\mathcalb{m}$, say between two curves $\pi$ and $\sigma$, never backtracks over vertices (only edges), so we already expect the error this introduces to be relatively small. Let $\mathcalb{m}^{+}$ denote the monotone morphing resulting from this simple strategy. Observe that
@@
We refer to this process of adding vertices to the two curves as refinement. In practice, in many cases, one or two rounds of (carefully implemented) refinement are enough to isolate the maximum leash in the morphing from the non-monotonicity, and followed by the above brute-force monotonization leads to the (practically) optimal Fréchet distance. Even for pathological examples, after a few more rounds of refinement, this process computes the exact Fréchet distance. That is, the computed lower bound, which is the VE-Fréchet distance, is equal to the width of the computed monotone morphing, which is the Fréchet distance.
Remark 1.6 (Floating point issues).
As we are implementing our algorithm using floating point arithmetic, and the calculation of the optimal Fréchet distance involves distances, impreciseness is unavoidable. A slight improvement in preciseness can be achieved by using squared distances (and also slightly faster code) — but for simplicity we have not used this idea in our code. In particular, we take the somewhat pragmatic view, that an approximation to the optimal up to a factor of (say) $1.00001$ can be considered as computing the “optimal” solution.
Note that Fréchet morphings are somewhat less sensitive to numerical issues than other geometric problems — indeed, once a morphing is computed, one can compute its width directly.

### v4

@@ Abstract
Combined with a Dijkstra/Prim type algorithm, this leads to a realization of the Fréchet distance (i.e., a morphing) that is locally optimal (aka locally correct), that is both easy to compute, and in practice, takes near linear time on many inputs. The new morphing has the property that the leash is always as short as possible. These matchings/morphings are more natural and are better than the ones computed by standard algorithms – in particular, they handle noise more graciously. This approach should make the Fréchet distance more useful for real-world applications.
We implemented the new algorithm and various strategies to obtain reasonably fast practical performance. We performed extensive experiments on our new algorithm, and released publicly available (and easily installable and usable) Julia and Python packages. Our algorithms can be used to compute the almost-exact Fréchet distance between polygonal curves.
Implementations and numerous examples are available here: frechet.xyz.
We emphasize, however, that the existing state-of-the-art algorithm/implementation in C++ is faster, by several orders of magnitude, than our current algorithm/implementation.
@@ 1.1 Definitions
Given two polygonal curves, their Fréchet distance is the length of a leash that a person needs if they walk along one of the curves. In contrast, a dog connected by the leash walks along the other curve, assuming they synchronize their walks to minimize the length of this leash. (I.e. they walk to minimize their maximum distance apart during the walk.) Our approach is slightly different than the standard approach, and we define it carefully first.
@@ 1.1.1 Free space diagram and morphings
The free space diagram $R$ is partitioned into a non-uniform grid, where each cell corresponds to all leash lengths when a point lies on a fixed segment of one curve, and the other lies on a fixed segment of the other curve, see Figure 1.1. For a given value $\delta\geq 0$, the sublevel set of a real-valued function consists of all inputs whose function value is $\leq\delta$. It is known that for any value $\delta\geq 0$, the sublevel set of the elevation function inside such a grid cell is a clipped ellipse.
@@
A morphing [^1] A morphing induces a natural homotopy between the two curves. $\mathcalb{m}$ between $\pi$ and $\sigma$ is a (not self-intersecting) curve $\mathcalb{m}\subseteq R\left(\pi,\sigma\right)$ with endpoints $(0,0)$ and $(\left\|{\pi}\right\|,\left\|{\sigma}\right\|)$. The set of all morphings between $\pi$ and $\sigma$ is $\mathcal{M}_{\pi,\sigma}$. A morphing that is a segment inside each cell of the free space diagram that it visits, is well behaved [^2] All the morphings we deal with in this paper are well behaved. .
@@ 1.2 Background
Alt and Godau [AG95] presented a rather involved $O(n^{2}\log n)$ time algorithm to compute the Fréchet distance using parametric search. The parametric search can be removed by using randomization, giving a simpler algorithm as shown in [HR14]. Buchin et al. [BBL+16] presented an alternative algorithm for computing the Fréchet distance that replaces the decision procedure by using a data structure to maintain appropriate lower envelopes. Despite some simplifications, all these algorithms are somewhat involved.
Unfortunately, it is believed this problem requires quadratic time in the worst case, although a logarithmic speedup is possible; see [BBMM17] and references therein. The quadratic time can be improved for realistic inputs by assuming that the input is “nice”, and introducing approximation, but the resulting algorithms are still not simple [DHW12].
@@ 1.2.1 Variants of the Fréchet distance
The weak Fréchet distance allows morphings where the agents are allowed to go back and traverse portions of the curve visited previously (i.e., the morphing does not have to be $x/y$-monotone). Since the weak Fréchet distance allows considering more parameterizations, it is potentially smaller, and we have that $\mathsf{d}^{w}_{\mathcal{F}}\left(\pi,\sigma\right)\leq\mathsf{d}_{\mathcal{F}}\left(\pi,\sigma\right)$, for any two curves $\pi,\sigma$. Elegantly, Alt and Godau [AG95] showed that the weak Fréchet distance can be reduced to computing the minimum spanning tree of an appropriate graph. Unfortunately, there does not seem to be a natural way to overcome this non-monotonicity (and thus get the “strong” version).
@@
The complexity of these algorithms, together with the sensitivity of the Fréchet distance to noise, leads to using “easier” related measures, such as the discrete version of the problem, and dynamic time-warping (discussed below). In the discrete version, you are given two sequences of points $p_{1},\ldots,p_{n}$, and $q_{1},\ldots,q_{m}$, and the purpose is for two “frogs” starting at $p_{1}$ and $q_{1}$, respectively, to jump through the points in the sequence until reaching $p_{n},q_{m}$, respectively, while minimizing the maximum distance between the two frogs during this traversal. At each step, only one of the frogs can jump from its current location to the next point in its sequence (no jumping back). Computing the optimal distance under this measure can be done by dynamic programming, similar to the standard approach to edit distance. Indeed, the configuration space here is the grid $H=\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$, where $\left\llbracket n\right\rrbracket=\{1,\ldots,n\}$.
To use the discrete version in the continuous case, one sprinkles enough points along both input curves and then solves the discrete version of the problem. Beyond the error this introduces, to get a distance that is close to the standard Fréchet distance, one has to sample the two curves quite densely in some cases.
For the (monotone) discrete Fréchet distance, the induced graph on the grid $H$ is a DAG, and the task of computing the Fréchet distance is to find a minimum bottleneck path from $(1,1)$ to $(n,m)$, where the weights are on the vertices. Here, the weight on the vertex $(i,j)$ is the distance $\left\|{p_{i}-q_{j}}\right\|$. In particular, a Fréchet morphing is an $x/y$-monotone path in $H$ from $(1,1)$ to $(n,m)$. The standard algorithm to do this traverses the grid, say, by increasing rows $i$, and in each row by increasing column $j$, such that the value at $(i,j)$ is the maximum of the length of the leash of this configuration, together with the minimum solution for (i1,j)(i-1,j) and (i,j1)(i,j-1). This algorithm leads to a straightforward, $O(nm)$ time algorithm for the discrete Fréchet distance. However, the morphing computed might be inferior, see Figure 1.2 for such a bad example.
@@
For simplicity, assume that the pairwise distances between all pairs of points in the two sequences are unique. We would like to imagine that we have a retractable leash that can become shorter at times, and the leash “resists” being longer than necessary. It is thus natural to ask for a morphing where the leash is as short as possible at any point in time.
Informally, the optimal retractable Fréchet morphing between the two sequences includes the bottleneck configuration, realizing the Fréchet distance, in the middle of its path, and the two subpaths from the endpoints to this configuration have to be also recursively optimal. This concept is formally defined and described in Section 3.2. This concept was introduced by Buchin et al. [BBMS19]. Interestingly, they show that the discrete version can be computed in $O(nm)$ time, but unfortunately, the algorithm is quite complicated. They also show that the continuous retractable Fréchet can be computed in $O(n^{3}\log n)$ time.
Buchin et al. [BBMS19] refers to this Fréchet distance as locally correct, but we prefer the retractable labeling. The term “retractable Fréchet” was used by Buchin et al. [BBL+16], but in a different (and not formally defined) context than ours.
@@
One way to get less sensitivity for noise is to compute the total area “swept” by the leash as the walk is being performed. In the discrete case, we add up the lengths of the leashes during the configurations in the walk. There is also work on extending this to the continuous setting [MP99, BBK+20]. For the continuous case, this intuitively boils down to computing (or approximating) an integral along the morphing. See Section 5.5 for more details.
@@ 1.2.2 Critical events
    Vertex-edge events: The distance between a vertex of one curve and an edge of the other.
@@ 1.2.3 Algorithm engineering the Fréchet distance
Given the asymptotic complexity and involved implementations of the aforementioned algorithms, there has been substantial work on practical aspects of computing the Fréchet distance. In particular, in 2017, ACM SIGSPATIAL GIS Cup had a programming challenge to implement algorithms for computing the Fréchet distance. See [WO18] for details.
More recently, Bringmann et al. [BKN21] presented an optimized implementation of the decider for the Fréchet distance. Somewhat informally, Bringmann et al. [BKN21] builds a decider for the Fréchet distance using a $kd$-tree over the free space diagram, keeping track of the reachable regions on the boundary of each cell, refining cells by continuing down the (virtual) $kd$-tree if needed.
@@ 1.3.1 Result I: A new algorithm for retractable discrete Fréchet
We observe that a natural approach to compute the retractable Fréchet morphing is to modify Dijkstra’s/Prim’s algorithm so that it solves the minimum bottleneck path problem. This observation leads to a simpler (but log factor slower) algorithm for computing it. The only modification of Dijkstra necessary is that one always handles the cheapest edge coming out of the current cut induced by the set of vertices already handled. (In the discrete Fréchet case, the weights are on the vertices, but this is a minor issue.) This modified version of Dijkstra is well known. Still, we include, for the sake of completeness, the proof showing that it indeed computes the recursively optimal path, which is also a retractable Fréchet morphing between the two sequences. This algorithm results in better and more natural morphings, see Figure 1.2.
@@
Figure 1.3: Two curves, and their associated VE-Fréchet graph. Note that every internal edge of the grid contains a portal (i.e., a vertex of the graph), but in many cases the portals of two edges are co-located on their common vertex (see the square-marked vertex in the middle figure). For the edges adjacent to the starting and ending corners, we set their portals to lie at the corners themselves. See here for more info about a similar example.
@@ 1.3.2 Result II: A new distance and algorithm: VE-Fréchet
Because of the continuity and strict convexity of the elevation function, the function has a unique minimum on each edge of the free space diagram grid – geometrically, this is the minimum distance between a vertex of one curve, and an edge of the other curve (a vertex-edge event). We restrict our solution to enter and leave a cell only through these portals (which are easy to compute). The continuous configuration space now collapses to a discrete graph that is somewhat similar to the natural grid graph on $\left\llbracket n\right\rrbracket\times\left\llbracket m\right\rrbracket$. Indeed, a grid cell has four portals on its boundary edges. Specifically, there are directed edges from the portal on the bottom edge to the portal on the top and right edges of the cell. Similarly, there are edges from the portal on the left edge to the portals lying on the top and right edges. As before, the values are on the portals, and our purpose is to compute the optimal bottleneck path in this grid-like graph. Examples of this graph are depicted in Figure 1.3, Figure 4.1 and Figure 4.2.
@@
Munich and Perona [MP99] used a similar idea, but they used it in the other direction – namely, in defining a better CDTW distance for two discrete sequences. However, this idea was already present (implicitly) in the original work of Alt and Godau [AG95] – indeed, their algorithm for the Weak Fréchet distance uses only the Vertex-Edge events (i.e., edges in the free space diagram). This problem boils down to solving the bottleneck shortest path problem in an undirected graph. In this case, this problem can be solved by computing the minimum spanning tree (e.g., by Prim’s algorithm, which is a variant of Dijkstra’s algorithm), as Alt and Godau do. For the directed case, one needs to use a variant of Dijkstra’s algorithm [GT88] – see Lemma 3.2. See also Buchin et al. [BBD+17] who also used a similar idea.
We can now run the retractable bottleneck shortest-path algorithm (i.e., the variant of Dijkstra described above) on this implicitly defined graph, computing the vertices and edges of it, as they are being explored. For many natural inputs, this algorithm does not explore a significant fraction of the configuration space, as it involves distances that are significantly larger than the maximum leash length needed. The algorithm seems to have near-linear running time for many natural inputs. The VE-Fréchet morphing is the one induced by the computed path in this graph. Unfortunately, the VE-Fréchet morphing might allow the agents to move backwards on an edge, but importantly, the motion across a vertex is monotone. Namely, the VE-Fréchet is monotone for vertices, but not necessarily monotone on the edges. A vertex is thus a checkpoint that once passed, cannot be crossed back.
@@ 1.3.3 Result III: New algorithm for the regular Fréchet distance
One can, of course, turn any morphing into a monotone one by staying put instead of moving back. This approach is appealing for VE-Fréchet, as the corresponding VE morphing $\mathcalb{m}$, say between two curves $\pi$ and $\sigma$, never backtracks over vertices (only edges), so we already expect the error this introduces to be relatively small. Let $\mathcalb{m}^{+}$ denote the monotone morphing resulting from this simple strategy. Observe that
@@
We refer to this process of adding vertices to the two curves as refinement. (See Figure 1.4.) In practice, in many cases, one or two rounds of (carefully implemented) refinement are enough to isolate the maximum leash in the morphing from the non-monotonicity, and followed by the above brute-force monotonization leads to the (practically) optimal Fréchet distance. Even for pathological examples, after a few more rounds of refinement, this process computes the almost-exact Fréchet distance. That is, the computed lower bound, which is the VE-Fréchet distance, is equal to the width of the computed monotone morphing, which is the Fréchet distance.
Remark 1.6 (Almost-exact: Floating point issues).
As we are implementing our algorithm using floating point arithmetic, and the calculation of the optimal Fréchet distance involves distances, imprecision is unavoidable. A slight improvement in precision can be achieved by using squared distances (and also slightly faster code) — but for simplicity, we have not used this idea in our code. In particular, we take the somewhat pragmatic view that an approximation to the optimal up to a factor of (say) $1.00001$ can be considered as computing the “optimal” solution. We refer to such solutions as being almost-exact.
Note that Fréchet morphings are somewhat less sensitive to numerical issues than other geometric problems — indeed, once a morphing is computed, one can calculate its width directly.
Remark 1.7 (What if one wants the exact distance?).
As pointed out above, our algorithm computes the almost-exact Fréchet distance, and in practice, there is no difference to the exact Fréchet distance (and in many cases they seem to coincide). Nevertheless, what if one insists on the exact Fréchet distance?
The refinement process can be modified to compute the monotonicity events on the regions on the curves where monotonicity is being violated. This change would readily lead to an exact algorithm, though we did not pursue it any further. See the paper [CHR25], who gave an exact implementation of our algorithm with some additional improvements.